# Tokenization is Sensitive to Language Variation

### v1

@@ 4.2 Tasks Sensitive to Language Variation
Second, we evaluate tokenizers on their downstream performance on tasks that are sensitive to language variation. The gold label for the task should use stylistic or form-based signals.
AV. Past work has found authorship verification (i.e., whether two texts were written by the same author) highly sensitive to tokenization, with significant gaps between the BERT and RoBERTa tokenizers (Zhu and Jurgens, 2021). Therefore, we curate a new authorship verification dataset of 40.8k train, 2.5k dev and 4.8k test pairs of texts from different domains, similar in distribution to the Miscellaneous corpus (c.f. §3). Labels are balanced in the test set. See details in Appendix LABEL:app:sadiri.
@@
NUCLE. We use the NUCLE 3.3 corpus Dahlmeier et al. (2013) for multi-label classification of the errors that were made by English learners in a given sentence. NUCLE was annotated by professional English instructors for 27 error types (e.g., verb tense or article use). It includes 22k unique sentences with errors. We add a sample of 5k error-free sentences from the same dataset. We split it into a train (80%) and dev (10%) set.
CORE. We use the Corpus of Online Registers of English Laippala et al. (2023) for register classification. Register is one of the most important factors associated with linguistic variation Biber (2012). We use 8 main register labels (e.g., spoken or informational description) for multi-class prediction. To increase the occurence of rarer labels, we split long texts and reach a train size of 30k and a dev size of 5k. See Appendix LABEL:app:sadiri for details.
Dialect. We randomly sample 60k instances from GLUE-dialect (§4.1) and the original GLUE task, to create a dialect classification task for five dialects and Standard American English in the original GLUE texts. We use 50k texts for the train and 5k for the dev set.
5 Modeling
For each investigated tokenizer (cf. §3), we pre-train one BERT model. We use encoder instead of decoder models, as encoder models tend to reach higher performance for classification tasks for low parameter settings. This allows us to train models using fewer GPU hours.
Pretraining BERT Models. We experimented with pre-training BERT models using a token count T that was exponentially bigger than the number of parameters P, similar to the original BERT papers Devlin et al. (2019); Turc et al. (2019). However, we find that for the same compute, using a bigger model size P and less tokens T improves the training loss. Chinchilla’s scaling law might also hold for smaller encoder models, specifically optimal parameter count could scale with the token size for a fixed compute $P_{OPT}\approx T^{23/27}$ Hoffmann et al. (2024). For the remainder of this work, we use the base BERT model architecture with 110M parameters, initialize all weights randomly and pre-train on 750M tokens sampled in sequences of 512 from the Miscellaneous corpus (§3.3) and use a batch size of 32 and 45k steps. For further details and hyperparameters, see Appendix LABEL:app:pretraining.
Fine-Tuning BERT. Unless otherwise specified, we used 3 epochs, a max sequence length of 128, a batch size of 32 and a learning rate of 2e-5 to perform the classification tasks. We evaluate on the dev set for GLUE. For comparability, we use the same setup for tasks requiring sensitivity to language variation. Only for the authorship verification task we use a contrastive training setup, then use the dev set to find an optimal cosine similarity threshold and calculate accuracy on the test set.
Figure 2: Pairwise significance testing of models. We use pairwise McNemar (1947)’s test to test how differently models trained with different tokenizers classify the tasks requiring robustness to language variation (first row) and tasks requiring sensitivity to language variation (second row). Tokenizers are sorted by mean performance. Blue colors show statistical significance, while red colors are above the 0.05 threshold.
6 Results
We show the performance of fine-tuned BERT models on tasks that require robustness to language variation in Table 2 and tasks that require sensitivity to language variation in Table 3. Note that the performance differences tend to be relatively small, which is consistent with previous work comparing different tokenizer algorithms on downstream tasks Ali et al. (2024); Schmidt et al. (2024). We compute the pairwise McNemar, 1947’s test for the pre-tokenizer, fitting corpus and vocabulary size settings, see Figure 2. We use the Bonferroni correction Dunn (1961) for our total of 26 tests.
RQ1: Tokenizer settings perform differently on tasks robust and sensitive to language variation. Tasks sensitive to language variation profit from tokenizers that encode more variation through: (i) a larger vocabulary size (§6.3) and (ii) pre-tokenizers that combine different Unicode Character Categories like punctuation and letters (§6.1).
Model orig +typo +dialect AVG
Corpus PMed 80.8 68.9 78.6 76.1
Wiki 80.7 68.3 79.3 76.1
Twitter 81.1 68.7 78.7 76.2
Pre-Tok
@@ 6 Results
81.7 68.4 79.7 76.6
Size 500 78.5 71.6 76.6 75.6
4k 80.1 69.8 78.2 76.0
32k 81.7 68.4 79.7 76.6
64k 81.2 68.3 79.4 76.3
128k 81.4 67.4 79.6 76.1
Table 2: Performance on Tasks Robust to Language Variation. We display BERT performances, averaged on the original four GLUE tasks and their perturbations using spelling mistakes (+typo) and dialectal transformations (+dialect). We boldface the best performances. For the averaging column (AVG), italics indicate tokenizers with no statistically significant difference from the best-performing tokenizer.
Model AV (acc) $\uparrow$ PAN (acc) $\uparrow$ CORE (acc) $\uparrow$ NUCLE (F1) $\uparrow$ DIALECT (F1) $\uparrow$ Mean
Corpus PubMed 81.7 65.6 56.5 23.0 88.0 63.0
Wikipedia 81.8 64.8 54.9 23.7 88.8 62.8
Twitter 83.4 66.6 56.7 22.1 88.1 63.4
@@
82.7 64.9 57.9 21.7 88.8 63.2
Size 500 79.2 62.7 51.6 14.1 58.7 58.7
4k 81.9 67.8 56.2 15.2 88.2 61.9
32k 82.7 64.9 57.9 21.7 88.8 63.2
64k 82.8 66.3 56.4 23.6 88.8 63.6
128k 82.7 67.1 54.3 22.8 88.7 63.1
Table 3: Performance on Tasks Sensitive to Language Variation. We display the BERT performance on our Authorship Verification (AV), the PAN, the CORE, and the multi-Dialect dataset. According to mean performance values, the best fitting corpus is Twitter, the best pre-tokenizer is
GPT
2 and the best model size is 64k. For the mean column, italics indicate tokenizers with no statistically significant difference from the best-performing tokenizer.
@@ 6.1 Pre-Tokenizers
Pre-tokenizers have the greatest influence on performance. For both task types, the range of performance values is largest for pre-tokenizers, second largest for vocabulary size and smallest for fitting corpus. This is surprising as, to the best of our knowledge, pre-tokenizers have received the least attention in previous work.
Pre-tokenizers improve performance. Using pre-tokenizers improves performance over using no pre-tokenizer. Combining leading whitespaces with letters (_ W S ) generally improved performance over isolating whitespaces ( W S ). A reason could be that splitting off whitespaces increases, in the worst case doubles, the number of input tokens to the model forcing it to reason over longer sequences Radford et al. (2019).
Combining Unicode Character Categories is more useful for tasks requiring sensitivity to language variation. For both task types, L L A M A 3 and G P T 2 are among the best performing pre-tokenizers. One of their main differences to other pre-tokenizers is that the combination of different Unicode Character Categories (e.g., punctuation and letters) within the same token is less often allowed. Overall, G P T 2 allows for the fewest combinations of Unicode Character Categories. By preventing combinations of categories like punctuation and letters, the vocabulary of G P T 2 can include a broader range of tokens that only consist of letters, e.g., “_queer” in Table 4. This could explain the better performance of G P T 2 on tasks that require robustness to language variation. In contrast, L L A M A 3 performed better on tasks requiring sensitivity to language variation. Allowing the mixing of one initial punctuation mark with letters, e.g., including “’all” in Table 4, seems to help L L A M A 3 solve PAN.
@@ 6.2 Fitting Corpus
Twitter performs best on tasks requiring sensitivity to language variation. This aligns with our expectation that the Twitter corpus include more spelling variation and a larger set of language varieties than PubMed and Wikipedia.
Twitter performs surprisingly well on tasks requiring robustness to language variation. Interestingly, Twitter performs indistinguishably from other fitting corpora on tasks that require robustness to language variation. Originally, we expected a more standardized corpus like Wikipedia to perform better, as it should lead to less spelling variation and thus more “full words” in the vocabulary (e.g., “precursor” in Table 4).
@@ 6.3 Vocabulary Size
A larger vocabulary size might be useful for tasks requiring sensitivity to language variation. Performance peaked at 32k for tasks requiring robustness to language variation and at 64k for tasks requiring sensitivity to language variation. It seems that a vocabulary size of 32k is sufficient to learn common standard forms and more tokens are needed to include style and form variations. For example, “jumper” is only included with vocabulary size 64k, see Table 4. Note that optimal vocabulary sizes might scale with model size Tao et al. (2024).
Smaller vocabulary is more robust to spelling variation. The vocabulary size of 500 consistently performs worse than larger vocabulary sizes. It seems that BERT struggles more with long input sequences and learning to compose words for small vocabulary sizes. An exception is the GLUE-Typo task, where the tokenizer of size 500 performs best. This is consistent with character-based tokenizers being more robust to spelling variation (cf. §3.1).
@@ 7 Pre-evaluating Downstream Tokenizer Impact
An popular alternative to training several LLMs to evaluate tokenizers is using fast intrinsic measures on the token distribution on the downstream task corpus. However, common intrinsic measures are not consistently correlated with downstream performance Rust et al. (2021); Zouhar et al. (2023); Schmidt et al. (2024). Further, they do not make use of task labels and can thus be considered task agnostic. Imagine a corpus on which you want to perform both, a task requiring robustness to language variation and a task requiring sensitivity to language variation (cf. Figure 1). Task-agnostic intrinsic measures like Rényi efficiency and Corpus Token Count will always predict the same performance for both tasks. However, we found that the type of task can have an influence on what tokenizer settings work better (cf. §6).
We experiment with a task-dependent measure: logistic regression with task labels as the dependent variable. Specifically, we use a tokenizer’s vocabulary as the feature set with a bag-of-token approach. For tasks with two texts as inputs, token combinations between sentence pairs serve as features. For multi-class and multi-label tasks, we use one-vs-rest models. See Appendix LABEL:app:log-regression for details.
@@
3). Note that allowing less Unicode Character Category combinations (e.g., numbers and letters) increases the broadness of the vocabulary (e.g.,
@@
2 allows for less combinations than other pre-tokenizers and is the only one including “queer” in its vocabulary).
@@ 7.1 Results
We display the Pearson correlation of Rényi efficiency, Corpus Token Count and logistic regression performances with the fine-tuned BERT performances in Table 5. Similar to Schmidt et al. (2024) we only find a light negative correlation with Rényi efficiency.
Corpus token count is conversely correlated with downstream performance, showing a negative correlation for tasks requiring robustness and a positive correlation for tasks requiring sensitivity to language variation. Corpus Token Count and correlated measures are often used to assess the compression quality of a tokenizer on a reference corpus or downstream task. A higher value entails more tokens in the corpus and thus worse compression. Better compression is commonly believed to be a sign of a better tokenizer, and has often been thought to be correlated with better downstream performance Rust et al. (2021); Ali et al. (2024); Velayuthan and Sarveswaran (2025). We show that the correlation is flipped for tasks that are robust and sensitive to language variation highlighting the difference in tokenizer requirements for the two task types.

### v2

@@ 4.2 Tasks Sensitive to Language Variation
Second, we evaluate tokenizers on tasks that are sensitive to language variation. The gold label for such tasks should be sensitive to stylistic or form-based signals.
AV. Models performing authorship verification (i.e., are two texts written by the same author?) usually need to be sensitive to the different styles and forms used by different authors (Zhu and Jurgens, 2021; Wegmann et al., 2022; Wang et al., 2023). Past work has found authorship verification to be sensitive to tokenization, with significant gaps between the BERT and RoBERTa tokenizers (Zhu and Jurgens, 2021). Therefore, we curate a new authorship verification dataset of 40.8k train, 2.5k dev and 4.8k test pairs of texts from different domains, similar in distribution to the Miscellaneous corpus (cf. §3). Labels are balanced in the test set. See details in Appendix LABEL:app:sadiri.
@@
NUCLE. We use the NUCLE 3.3 corpus Dahlmeier et al. (2013) for multi-label classification of the errors that were made by English learners in a given sentence. NUCLE was annotated by professional English instructors for 27 error types (e.g., verb tense or article use). From NUCLE, we take all 22k unique sentences containing errors and sample 5k error-free sentences. We split it into a train (80%) and dev (10%) set.
CORE. We use the Corpus of Online Registers of English Laippala et al. (2023) for register classification. Register is one of the most important factors associated with linguistic variation Biber (2012). We use the 8 main register labels (e.g., spoken or informational description) for multi-class prediction. To increase the occurence of rarer labels, we split long texts and reach a train size of 30k and a dev size of 5k. See Appendix LABEL:app:sadiri for details.
Dialect. We randomly sample 60k instances from GLUE+dialect (§4.1) and the original GLUE task, to create a dialect classification task with five dialects and Standard American English in the original GLUE texts. We use 50k texts for the train and 5k for the dev set.
Tokenizer Setting Model GLUE +typo +dialect AVG
Fitting Corpus PubMed 80.8 $\pm$ 0.0 69.1 $\pm$ 0.2 78.6 $\pm$ 0.2 76.2 $\pm$ 0.0
Wikipedia 80.7 $\pm$ 0.3 68.6 $\pm$ 0.2 79.3 $\pm$ 0.2 76.2 $\pm$ 0.2
Twitter 81.1 $\pm$ 0.0 69.1 $\pm$ 0.5 78.8 $\pm$ 0.1 76.4 $\pm$ 0.2
Pre-Tokenizer
@@
81.3 $\pm$ 0.4 68.2 $\pm$ 0.2 79.2 $\pm$ 0.4 76.2 $\pm$ 0.3
Vocabulary Size 500 77.2 $\pm$ 2.5 70.3 $\pm$ 2.6 75.6 $\pm$ 2.0 74.4 $\pm$ 2.4
4k 80.5 $\pm$ 0.8 70.3 $\pm$ 0.9 78.6 $\pm$ 0.8 76.4 $\pm$ 0.8
32k 81.3 $\pm$ 0.4 68.2 $\pm$ 0.2 79.2 $\pm$ 0.4 76.2 $\pm$ 0.3
64k 80.8 $\pm$ 0.4 67.6 $\pm$ 0.6 79.2 $\pm$ 0.2 75.9 $\pm$ 0.4
128k 78.7 $\pm$ 2.0 64.6 $\pm$ 1.9 76.1 $\pm$ 2.6 73.1 $\pm$ 2.2
Table 2: Performance on tasks requiring robustness to language variation. We display BERT performances, averaged on the original four GLUE tasks and their perturbations using spelling mistakes (+typo) and dialectal transformations (+dialect). We provide the mean and standard deviation ($\pm$) over three seeds respectively. We boldface the best performances for each column and investigated setting. For the averaging column (AVG), italics indicate tokenizers with no statistically significant difference from the best-performing tokenizer (cf. Figure 2).
5 Modeling
For each investigated tokenizer (cf. §3), we pre-train three BERT models with different seeds. We use encoder instead of decoder models, as encoder models tend to reach higher performance for classification tasks for low parameter settings. This allows us to train models using fewer GPU hours.
Pretraining BERT Models. We experiment with pre-training tiny BERT models using a token count T close to 3300M that is exponentially bigger than the 4.6M parameters P, similar to the original BERT papers Devlin et al. (2019); Turc et al. (2019). However, we find that for the same compute, using a bigger model size P and fewer tokens T improves the training loss. Chinchilla’s scaling law might also hold for smaller encoder models, specifically optimal parameter count could scale with the token size for a fixed compute $P_{OPT}\approx T^{23/27}$ Hoffmann et al. (2024). For the remainder of this work, we use the base BERT model architecture with 110M parameters, initialize all weights randomly and pre-train on 750M tokens sampled in sequences of 512 from the Miscellaneous corpus (§3.3) and use a batch size of 32 and 45k steps. For further details and hyperparameters, see Appendix LABEL:app:pretraining.
Fine-Tuning BERT. Unless otherwise specified, we use 3 epochs, a max sequence length of 128, a batch size of 32 and a learning rate of 2e-5 to perform the classification tasks. We evaluate on the dev set for GLUE. For comparability, we use the same setup for tasks requiring sensitivity to language variation. Only for the authorship verification task we use a contrastive training setup, then use the dev set to find an optimal cosine similarity threshold and calculate accuracy on the test set.
6 Results
We show the performance of fine-tuned BERT models on tasks that require robustness to language variation in Table 2 and tasks that require sensitivity to language variation in Table 3. When we investigate a specific setting (e.g., the fitting corpus in the first three rows in Table 2), we only change that setting and leave the other at their “default settings” to ensure comparability and isolate the effect of each individual setting without exhaustive testing of all possible combinations. We use the following default values for the three settings: the miscellaneous fitting corpus, the G P T 2 pre-tokenizer and a vocabulary size of 32k.
Note that the performance differences tend to be relatively small, which is consistent with previous work comparing different tokenizer algorithms on downstream tasks Ali et al. (2024); Schmidt et al. (2024). To ensure significance, we compute the pairwise McNemar, 1947’s test for the pre-tokenizer, fitting corpus and vocabulary size settings, see Figure 2. For the significance testing, we consider classifications by models with the same settings but different seeds to be stemming from the same rater. We use the Bonferroni correction Dunn (1961) for our total of 26 tests.
RQ1: Tokenizer settings perform differently on tasks robust and sensitive to language variation. Overall, tasks sensitive to language variation profit from tokenizers that encode more variation through a larger vocabulary size than 4k (§6.3). Note that the best-performing tokenizer settings are not always consistent across the individual variation-robust and variation-sensitive tasks (e.g., vocabulary size for GLUE+typo and GLUE in Table 2). We suspect that this is due to differences in the types of language variation present in the specific task datasets (e.g., character-level tokens seem to be more robust to spelling variation). Future work should investigate different types of language variation individually (e.g., spelling vs. lexical variation).
Model AV (acc) $\uparrow$ PAN (acc) $\uparrow$ CORE (acc) $\uparrow$ NUCLE (F1) $\uparrow$ Dialect (F1) $\uparrow$ AVG
Corpus PMed 81.7 $\pm$ 0.1 65.2 $\pm$ 0.5 55.9 $\pm$ 0.6 21.8 $\pm$ 1.2 87.9 $\pm$ 0.1 62.5 $\pm$ 0.4
Wiki 81.9 $\pm$ 0.1 65.5 $\pm$ 0.5 55.5 $\pm$ 0.6 23.5 $\pm$ 0.2 88.9 $\pm$ 0.5 63.1 $\pm$ 0.2
Twitter 82.9 $\pm$ 0.6 66.7 $\pm$ 0.4 56.5 $\pm$ 0.6 21.4 $\pm$ 0.9 88.3 $\pm$ 0.2 63.2 $\pm$ 0.2
@@ 6 Results
82.6 $\pm$ 0.5 66.6 $\pm$ 1.2 56.3 $\pm$ 1.2 21.8 $\pm$ 0.9 88.4 $\pm$ 0.6 63.1 $\pm$ 0.4
Size 500 78.2 $\pm$ 0.9 62.6 $\pm$ 0.2 51.1 $\pm$ 0.6 13.1 $\pm$ 0.9 85.6 $\pm$ 0.4 58.1 $\pm$ 0.6
4k 81.9 $\pm$ 0.1 67.8 $\pm$ 0.6 55.1 $\pm$ 1.0 17.4 $\pm$ 2.8 87.9 $\pm$ 0.7 62.0 $\pm$ 0.6
32k 82.6 $\pm$ 0.5 66.6 $\pm$ 1.2 56.3 $\pm$ 1.2 21.8 $\pm$ 0.9 88.4 $\pm$ 0.6 63.1 $\pm$ 0.4
64k 82.7 $\pm$ 0.4 67.2 $\pm$ 0.7 54.9 $\pm$ 1.5 22.0 $\pm$ 1.1 88.1 $\pm$ 0.6 63.0 $\pm$ 0.6
128k 80.1 $\pm$ 2.2 62.4 $\pm$ 3.4 51.2 $\pm$ 2.2 19.0 $\pm$ 2.7 81.5 $\pm$ 5.1 58.9 $\pm$ 3.0
Table 3: Performance on tasks requiring sensitivity to language variation. We display BERT performances on our Authorship Verification (AV), the PAN, the CORE, and the multi-Dialect dataset. We provide the mean and standard deviation ($\pm$) over three seeds respectively. For the averaging column (AVG), italics indicate tokenizers with no statistically significant difference from the best-performing tokenizer (cf. Figure 2).
@@ 6.1 Pre-Tokenizers
Pre-tokenizers have the greatest influence on performance. For both task types, the range of performance values is largest for pre-tokenizers, second largest for vocabulary size and smallest for fitting corpus. This is surprising as, to the best of our knowledge, pre-tokenizers have received the least attention in previous work. For both task types, using pre-tokenizers improves performance over using no pre-tokenizer.
Pre-tokenizer performance differs more between individual tasks than between task types. For both task types, _ W S is among the best performing pre-tokenizers. Combining leading whitespaces with letters (_ W S ) generally improved performance over isolating whitespaces ( W S ). Notably, _ws cuts the input tokens to a model in half as single whitespaces between words are not separate tokens anymore. For dialect and spelling-transformed GLUE, CORE, NUCLE and Dialect, the whitespace-based pre-tokenizers perform the best. For AV, PAN and the original GLUE tasks, G P T 2 and L L A M A perform better. Tasks like grammatical error detection (NUCLE) could be seen as identifying deviations from a standard, and might benefit from pre-tokenizers that include tokens with typical combinations of Unicode character categories. One of the main differences of G P T 2 and L L A M A to other pre-tokenizers is that the combination of different Unicode character categories (e.g., punctuation and letters) within the same token is less often allowed. By preventing combinations of categories like punctuation and letters, the vocabulary of G P T 2 can include a broader range of tokens that only consist of letters, e.g., “_queer” in Table 4. This could explain the better performance of G P T 2 on the original GLUE task. In contrast, L L A M A 3 also allows the mixing of one initial punctuation mark with letters, e.g., including “’all” in Table 4, which after manual inspection seems to especially help L L A M A 3 solve PAN. Future work could investigate different individual tasks and the influence of pre-tokenizers individually.
@@ 6.2 Fitting Corpus
On tasks requiring sensitivity to language variation, Twitter performs best. This aligns with our expectation that the Twitter corpus includes more spelling variation and a larger set of language varieties than PubMed and Wikipedia. Interestingly, Wikipedia performs the best on NUCLE and Dialect. Grammatical error detection (NUCLE) could be seen as identifying deviations from a standard, and might benefit from training on corpora with few grammatical errors — like Wikipedia.
On tasks requiring robustness to language variation, Twitter performs surprisingly well. Interestingly, Twitter performs indistinguishably from other fitting corpora on tasks that require robustness to language variation. Originally, we expected a more standardized corpus like Wikipedia to perform better on the original GLUE task, as it should lead to less spelling variation and thus more “full words” in the vocabulary (e.g., “precursor” in Table 4).
@@ 6.3 Vocabulary Size
A larger vocabulary size might be useful for tasks requiring sensitivity to language variation. Performance peaked at 4k for tasks requiring robustness to language variation and at 32k for tasks requiring sensitivity to language variation. It seems that a smaller vocabulary size is sufficient to learn common standard forms and more tokens are needed to include style and form variations. For example, “jumper” is only included with vocabulary size 64k, see Table 4. Note that optimal vocabulary sizes might scale with model size Tao et al. (2024). For both task types, the standard deviation is highest for the smallest (500) and the largest vocabulary size (128k).
Smaller vocabulary is more robust to spelling variation. The vocabulary size of 500 consistently performs worse than larger vocabulary sizes. It is possible that BERT struggles with long input sequences and learning to compose words for small vocabulary sizes. An exception is the GLUE+typo task, where the tokenizer of size 500 performs best. This is consistent with character-based tokenizers being more robust to spelling variation (cf. §3.1).
Figure 2: Pairwise significance testing of models. We use pairwise McNemar (1947)’s test to test whether there is a significant difference between classifications done by models trained with different tokenizers on the tasks requiring robustness to language variation (first row) and tasks requiring sensitivity to language variation (second row). Tokenizers are sorted by mean performance. Blue colors show statistical significance, while red colors are above the 0.05 threshold.
@@ 7 Pre-evaluating Downstream Tokenizer Impact
3). Note that allowing fewer Unicode character category combinations (e.g., numbers and letters) increases the broadness of the vocabulary (e.g., in our fitting corpora,
@@
2 allows for fewer combinations than other pre-tokenizers and is the only one to include “queer” in its vocabulary).
A popular alternative to training several larger LMs to evaluate tokenizers is using fast intrinsic measures on the token distribution on the downstream task corpus. However, common intrinsic measures are not consistently correlated with downstream performance Rust et al. (2021); Zouhar et al. (2023); Schmidt et al. (2024). Further, they do not make use of task labels and can thus be considered task agnostic. Imagine a corpus on which you want to perform both a task requiring robustness to language variation and a task requiring sensitivity to language variation (cf. Figure 1). Task-agnostic intrinsic measures like Corpus Token Count and Rényi efficiency Zouhar et al. (2023) will always predict the same performance for both tasks. However, we found that the type of task can have an influence on what tokenizer settings work better (cf. §6).
7.1 Logistic Regression Measure
We experiment with a task-dependent measure: the performance [^4] We use the same F1 and accuracy performance metrics as for BERT. of logistic regression with task labels as the dependent variable. Specifically, we use a tokenizer’s vocabulary as the feature set with a bag-of-token approach. For tasks with two texts as inputs, token combinations between sentence pairs serve as features. For multi-class and multi-label tasks, we use one-vs-rest models. See Appendix LABEL:app:log-regression for details.
@@ 7.2 Results
We display the Pearson correlation of Rényi efficiency, Corpus Token Count and logistic regression performances with the fine-tuned BERT performances in Table 5. Similar to Schmidt et al. (2024) we find a light negative correlation of LM performances with Rényi efficiency on NLU tasks.
Corpus token count has a negative correlation for tasks requiring robustness and a positive correlation for tasks requiring sensitivity to language variation. Corpus Token Count and correlated measures are often used to assess the compression quality of a tokenizer on a reference corpus or downstream task corpus. A higher value entails more tokens in the corpus and thus worse compression. Better compression is commonly believed to be a sign of a better tokenizer, and has often been thought to be correlated with better downstream performance Rust et al. (2021); Ali et al. (2024); Velayuthan and Sarveswaran (2025). We show that the correlation is flipped for tasks that are robust and sensitive to language variation highlighting the difference in tokenizer requirements for the two task types.